OpenStudy (anonymous):

Marco has a CD rack that contains 120 CDs arranged in no particular order. Fifteen CDs are pulled randomly from the rack. The results are three country, ten rap, and two alternative. Based on this information, what is the total number of country CDs you can expect to find in Marco's CD rack?

OpenStudy (anonymous):

15 CDs out of 120 CDs means you pulled 15/120 = 1/8 of the CDs. So you could multiply the 3 country CDs in the sample times 8 to get an expected number in the whole 120 CD set.

OpenStudy (anonymous):

Does that make sense? It of course doesn't guarantee you will find that number of country CDs... but if you only had the information from your sample, you might at least expect that the rest of the CDs had a similar spread of country vs. rap vs. alt.

OpenStudy (anonymous):

so I would get 21?

OpenStudy (anonymous):

I think you made an error :) if you found 3 country CDs out of 15, then you would expect to find 8 times as many in 120. So 8 x 3 = ?

OpenStudy (anonymous):

24 lol

OpenStudy (anonymous):

Another way to set this up is just as ratios... 3/15 = x/120 Then multiply both sides by 120 to get x.

OpenStudy (anonymous):

24 sounds better :)

OpenStudy (anonymous):

Thanks so much for helping me understand this better!

OpenStudy (anonymous):

Glad to help!